OpenStudy (anonymous):

The center of a circle is (h, 7) and the radius is 10. The circle passes through (3, -1). Find all possible values of h.

OpenStudy (raden):

(h - 3)^2 + (7 - (-1))^2 = 10^2 solve for h

OpenStudy (raden):

did u get it ?

OpenStudy (raden):

well, hope you understand this (h - 3)^2 + (7 - (-1))^2 = 10^2 (h - 3)^2 + 64 = 100 (h - 3)^2 = 100 - 64 (h - 3)^2 = 36 h-3 = +- sqrt(36) h-3 = +- 6 h = 3 +- 6 h1 = 3+6 = 9 h2 = 3-6 = -3 so, the possibles values of h = (-3, 9)